OpenStudy (anonymous):

4k-3=2k+5 4k-3-2k=2k+5-2k 2k-3=5 2k-3+3=5 2k=8 2k/2=8/2 k=4 but how do i justify it??????????????????

jimthompson5910 (jim_thompson5910):

Plug k = 4 back into the original equation and evaluate 4k-3=2k+5 4(4)-3=2(4)+5 16-3=8+5 13=13 Since the last equation is true, this verifies the answer.

OpenStudy (anonymous):

well the paper says like a. b. c. d. e. f. how do i put it into that

jimthompson5910 (jim_thompson5910):

well they expect 6 lines, but somehow I got the check in 4....hmm

jimthompson5910 (jim_thompson5910):

maybe something like a. 4k-3=2k+5 b. 4(4)-3=2(4)+5 c. 16-3=2(4)+5 d. 16-3=8+5 e. 16-3=8+5 f. 13=13

jimthompson5910 (jim_thompson5910):

does that help?

OpenStudy (anonymous):

hey what about 2(8u+2)=3(2u-7) 16u+4-6u-21 16u+4-6u=6u-21-6u 10u=4=(-21) 10u+4-4=(-21)-4 10u=-25 10u/10=-25/10 u=2.5

jimthompson5910 (jim_thompson5910):

Answer should be -2.5 or -5/2 Is that what you really meant?